OpenStudy (anonymous):

What is a good question for 8th graders struggling in Math right now?

OpenStudy (liv1234):

Do you need help?

OpenStudy (theraggedydoctor):

Google can help you.

OpenStudy (nuttyliaczar):

And we gotchu

OpenStudy (steve816):

A good question for them to try is... try to find a number that satisfies the following equation:

OpenStudy (steve816):

\[a^n + b^n = c^n\]

OpenStudy (steve816):

where n > 2

OpenStudy (anonymous):

Thanks! Just curious Steve816. Why did you choose that?

OpenStudy (steve816):

Because... that is a secret :)

OpenStudy (amistre64):

consider; what is a number?

OpenStudy (anonymous):

High school entrance exams, trying to design questions to practice

OpenStudy (steve816):

Oh, are you a teacher?

OpenStudy (amistre64):

i think that people who struggle with math are those who have tried to shortcut themselves by trying to memorize formulas and FOILage ... instead of learning what the fundamentals of math are.

OpenStudy (anonymous):

No, not a teacher, but trying to help a friend

OpenStudy (steve816):

@amistre64 I strongly agree, most of the time, the teachers tell you to do this and do that, and boom, it will magically work. But I am always intrigued by HOW it works, and it really bothers me if I do math problems without knowing how it is working.

OpenStudy (ikram002p):

Hmmm 8th grade kids need to know concept of rational and irrational also area of circle and ratios. And they also need to learn sets in basic way. Hmmm this all that I remember lol

OpenStudy (steve816):

Like for example, someone please explain how the formula for the volume of a sphere works, it just seems so random and intriguing because I don't know how people can figure out such things.

OpenStudy (amistre64):

volume is mostly a calculus modus operandi

OpenStudy (anonymous):

@amister64 definitely agree.

OpenStudy (steve816):

I don't even know what you're talking about, haven't gotten into calculus yet, but dang, can't wait till AP calc next year, seems so fascinating.

OpenStudy (ikram002p):

There's actually nice tricks and ways to do this, without calculus and they are fun as also they can make experiments hmmm

OpenStudy (bubblegum.):

try to work out a proof for the Pythagoras theorem

OpenStudy (amistre64):

\[2\pi\int^{r}_{0}(\sqrt{r^2-x^2})^2~dx\] \[2\pi\int^{r}_{0}r^2-x^2~dx\] \[2\pi(r^3-\frac13r^3)\] \[2\pi(\frac23r^3)\]

OpenStudy (mathmale):

@paulgimenez: I see that this question is your first post. Typically, students post actual math questions with which they'd like help. My first impulse was to delete your post, because it does not address a specific math problem that you'd like to solve. I'm leaving your post as is because several people have brought up interesting discussion points, but am closing it so others will not join. Welcome to OpenStudy; hope you find it to be ver profitable.

OpenStudy (mathmale):

Second thought: Are you an "8th grader struggling in math right now?" If so, just choose any problem you want from your homework and post that problem. We'll see where the discussion goes from there.

OpenStudy (amistre64):

As I see it, the question is math related, and therefore relevant to the subject. It needs not be deleted or closed by force.

OpenStudy (steve816):

I agree with amistre64, it indeed is math related, and there is no reason to close it.

OpenStudy (anonymous):

@mathmale, no, my cousin is presently in 8th grade and has high school entrance exams (in NYC you have to apply to high school). I don't know the questions I should be trying to help him with, so I searched for education social networks and open study came up. Figured it would be a good place to start.

OpenStudy (anonymous):

that's the specific reason for the question...

OpenStudy (amistre64):

if you need practice problems, a good site to use is interactmath.com enter, pick a textbook related to your course material, and they present you with ample practice problems and helpful solutions options.

OpenStudy (mathmale):

I see, and truly appreciate your clarifying this. Good of you to want to help your cousin. My first thought would be to recommend that you and your cousin check out the 'Net for "8th grade math curriculum" or just "8th grade math." Bet you'll find an interesting assortment of problems listed.

Nnesha (nnesha):

algebraic expression/equations how to translate sentence into math equations/expression they should know the meaning of each word in math form o^_^o so many students are struggling wih this topic. They should know how solve for x atleast and an exponent rules and also sign chart ;) -*-=+ +*+=+ -*+=-

OpenStudy (anonymous):

@amistre64 that's fantastic. Plus, I do like that the conversation expanded in directions I did not expect. Thanks gang.

OpenStudy (mathmale):

Paul: Here's one of my first search results: https://www.gomathacademy.com/pages/grade-8 This lists topics that would normally be addressed in 8th grade math.

OpenStudy (mathmale):

Perhaps your cousin would enjoy and benefit from OpenStudy. You could share this list of topics with him / her and suggest that he / she look for relevant problems, then work on them, asking here for help as needed.